### v1

# Towards Data-Driven Metrics for Social Robot Navigation Benchmarking

Abstract
This paper presents a joint effort towards the development of a data-driven Social Robot Navigation metric to facilitate benchmarking and policy optimization. We provide our motivations for our approach and describe our proposal for storing rated social navigation trajectory datasets. Following these guidelines, we compiled a dataset with 4427 trajectories—182 real and 4245 simulated—and presented it to human raters, yielding a total of 4402 rated trajectories after data quality assurance. We also trained an RNN-based baseline metric on the dataset and present quantitative and qualitative results. All data, software, and model weights are publicly available.
I Introduction
Social robot navigation is becoming an increasingly relevant field, with both growing academic research interest and notable industry involvement. Industrial applications include autonomous vehicles, robots for restaurants and museums, and delivery pods for last mile delivery [15].
Figure 1: We acquire trajectories from both real and simulated scenarios. Simulated trajectories have a lower cost and allow for ethical recording of unsafe behavior, but real trajectories are still required to ensure generalisation. We generate a top-down view for each trajectory and show it to raters to collect scores. We have made all the data we collected public, including trajectories, raters’ data, and their ratings. Metric and policy learning are use cases for the dataset.
Despite efforts of the social robot navigation community to establish principles and guidelines for evaluation [4], benchmarking of social robotics is not yet standardized and can be biased to suit specific research narratives. Social robot navigation (hereafter SocNav) is currently assessed through metrics focusing on success, safety, proxemics, or readability [10]. However, different research works consider different metrics depending on what their authors consider relevant and often make claims without considering tradeoffs between metrics or Pareto optimality for the metrics selected.
Furthermore, task type and navigation context [4] are often overlooked. How the type of task affects behavior is exemplified by robots entering people’s personal space, which is required if a robot is delivering an object to a person, but normally should be avoided. The importance of context is exemplified by whether a robot should pass bystanders at high speed, which is more appropriate when a robot is carrying urgent medical equipment than when returning to base to recharge, even though the task might be the same from a geometric perspective.
Many existing metrics are computed algorithmically from variables, or analytic [4]. Analytic metrics often have poor variable-wise scalability, and disregard factors such as the type of task or contextual factors such as the distribution of objects around the robot. Analytic metrics also frequently lack experimental support showing that what the metric measures matches user perceptions of robot behavior quality. Nevertheless, new metrics continue to be proposed, and, although the intuitions behind them are generally solid, they also often lack empirical support. We argue that metrics should be empirically validated, ideally learned from data, and should evaluate the overall effectiveness and socialness of an entire trajectory. We call such a metric an All-encompassing Learned Trajectory-wise (ALT) metric [4]. An ALT metric is learned directly from human evaluations of full robot trajectories, conditioned on context—a category of metrics which, to our knowledge, does not yet exist.
Motivated by the lack of a standardized method to benchmark SocNav policies, limitations of existing analytic metrics, and the absence of an ALT metric, this article studies the feasibility of a data-driven solution to create an ALT metric, which would solve all these problems.
Robots are not yet commonly integrated into everyday human environments with explicit quantitative feedback [20], so few datasets suitable for training learned metrics exist. This gap makes compiling a dataset specifically for this purpose essential. Therefore, this paper contributes with:
  1. 1.
    a specification for storing social robot navigation trajectories along with their ratings,
  2. 2.
    open-source tools for visualizing this data,
  3. 3.
    a proof-of-concept dataset following this spec, and
  4. 4.
    a learned metric exploiting this dataset.
Our dataset specification includes information concerning the environment, the humans, the robot’s task, its context, and the trajectory of the robot, as well as demographic information about the human raters and the scores for the trajectories they rated. The specification and any dataset that follows it can be exploited for (see Fig.1):
  • Training ALT metrics: Given a sufficiently large and varied set of trajectory-score pairs, a supervised machine learning model can be trained on the dataset to produce a learned ALT metric.
  • Policy development: Given a high-quality ALT metric, SocNav policies can be derived by optimizing the score for the metric—for example, using reinforcement learning or model predictive control.
Our proof-of-concept metric shows that an ALT metric can be learned, and encourages the community to help scale the dataset to the point where metrics learned over it can be used as the de facto metric for benchmarking robot trajectories [^1] Instructions to contribute are in the software repository https://github.com/SocNavData/SocNav3.. Policy development is out of the scope of this paper.
In addition to our main contributions, we provide experimental results on data quality and the accuracy of our metric, as well as qualitative results for stereotypical scenarios.
II Related Work
Regarding learned SocNav metrics, SocNav1 [8] and SocNav2 [2] are the initiatives that most closely resemble the goals of the paper at hand. The SocNav1 [8] dataset provides scores of static scenarios including one robot and a variable number of humans, with generic object bounding boxes and explicit interactions (human-to-human and human-to-object). The SocNav2 dataset [2] extended SocNav1 to account for brief periods of time and specific goals, but not full trajectories. Even though these works provide learned SocNav metrics, they are step-wise, i.e., they account for snapshots rather than whole trajectories. Furthermore, none of them consider the context of the tasks.
Although not strictly a metric, another notable research direction estimates the perturbation of pedestrian movement as a variable to minimize  [5, 1]. Despite its elegant simplicity, this approach is dependent on the number of people in the environment, and does not account for the context of the task nor the general perception of trajectory quality. Similarly, [13] learns a cost function over observations of real human motion interactions found in publicly available datasets, then uses that cost to determine a navigation controller. While this cost function could be used as a metric, this approach only considers the context of crossing pedestrians on a pavement and has the same shortcomings as [5] and [1].
Many datasets relevant to human trajectory modeling exist, but datasets which also include humans, robots and environmental data (such as a map or objects) are scarcer. JRDB [9], THOR [14], SCAND [7] and SACSoN [5] are examples of such datasets. They provide raw sensor data from teleoperated or autonomous robot trajectories, and frequently include a video feed and 3D LiDAR or RGBD data. However, with the exception of THOR [14], pedestrians are annotated as bounding boxes or 2D points without orientation.
We found no datasets containing rated robot trajectories, so we chose for this paper to reuse and annotate existing datasets along with recording new additional data. Although THOR initially seemed a good fit for this purpose because it includes object annotations, oriented pedestrians and even human gaze orientation, we chose SACSON instead because of its scale. Nevertheless, as mentioned in the introduction, we combined synthetic data with this real data so we could vary the quality of the trajectories in order to reduce the bias of the learned metric.
III Dataset Specification and Tools
This section first describes our proposed dataset specification, including the variables it represents and its file format, and then summarizes the software we developed to work with compatible datasets. The design decisions behind this specification are the result of a series of focus group sessions where the authors of the paper participated.
III-A Variables
The dataset comprises variables related to raters, trajectories, and rater-trajectory scores. For every rater, the dataset contains their age (integer), gender (categorical: female, male, non-binary, transgender, other, no-answer), country of residence (categorical), and a rating list. Each rating list contains tuples $(t,c,r)$, where $t$ is a trajectory identifier (string), $c$ is a context (string), and a $r$ is a score (float, $r\in[0,1]\subset\mathbb{R}$). We chose to acquire real-valued scores rather than a Likert scale following best-practice HRI studies [18] which regard sensitivity as more relevant than repeatability. To score trajectories, the survey tool used a slider with three references “extremely bad” (score 0), fair (score $0.5$) and “extremely good” (score $1$). Note that contexts are not necessarily bound to trajectories when trajectories are recorded; this enables us to vary the context to explore how different contexts affect the perception of the same trajectory.
Trajectories contain data on the robot, the task and its context, humans, objects, and the environment, as shown in Table I. Except for variables related to the task and the environment, which apply to the whole trajectory, variables are recorded with a timestamp at each time step.
The robot is described providing its pose (in $SE(2)$), speed (a twist vector), drive system (categorical variable, “differential”, “omnidirectional”, “ackerman”, or “biomimetic”) and shape (circle, rectangle, or a generic polygon described using a vertex representation). The focus group considered whether to add 3D information about the shape of robots or objects. While this could help detect potentially problematic situations (e.g., robot or object geometry impeding the visibility of other agents), most current metrics are very far from considering such aspects. Therefore, the focus group decided not to include 3D shapes in the initial version of the specification.
The goal is defined differently depending on the task type, which is provided as a categorical variable “go-to”, “guide-to”, “follow”, or “interact-with”. For go-to tasks, the specification requires a target 2D position (in meters) and a target angle (in radians), both with thresholds using Euclidean and angular distance, respectively. By varying these threshold values, a go-to task can specify a very stringent $SE(2)$ point target, just a position, or only an angle. Tasks of type guide-to, are specified similarly to go-to tasks, but require an additional human identifier, and the task target applies to the human, not the robot. Tasks of type follow and interact-with require a human identifier to specify which human is the robot expected to follow, along with an Euclidean distance threshold.
Humans are described with a numerical identifier, their 2D pose and, optionally, the 3D coordinates of their COCO-18 key-points. The focus group decided making the 3D version of the human pose mandatory would limit adoption, as few research groups or even companies have access to full human pose trackers. However, the focus group decided to add 3D human pose as an optional element so the specification would not become prematurely obsolete.
Objects are described with an identifier, a text-based description, their 2D pose and their shape (circle, rectangle, or polygon). The focus group chose a text-based description over a predefined set of object types because of its flexibility (text enables open-set objects) and fidelity (language models can often easily turn text into categories, but the reverse is not true without loss of information).
The remaining environmental information consists of the walls (specified as 2D polylines), a 2D grid map (with the same fields as ROS2 maps), and area semantics (text-based, following the same rationale as the text-based object description). The grid map is specified with a 2D data array, and the resolution of the grid is specified in metres per cell.
Type Variable Description
Pose 2D position (m) and orientation (rad) on the plane.
Speed Lineal (m/s), angular (rad/s)
Drive Categorical (differential/omni/ackerman)
Robot Shape 2D circle (radius), rectangle (width, height) or polygon (list of points).
Type Task type, either “go-to", “guide-to" or “follow".
Position + threshold For go-to and guide-to types of tasks. 2D position + threshold (m).
Orientation + threshold For go-to and guide-to types of tasks. 2D orientation + threshold (rad).
Human identifier For guide-to and follow types of tasks.
Task Context Textual description of the context, in English.
Identifier Integer that uniquely identifies the human in a given episode.
Pose 2D position (m) and orientation (rad) on the plane.
Humans Full Pose (optional) The 3D position of the COCO-18 key point set.
Identifier Integer that uniquely identifies the object in a given episode.
Type Free text describing the type of the object.
Pose 2D position (m) and orientation (rad) on the plane.
Objects Shape The 2D shape of the object.
Walls Sequence of polylines (2D, m).
Grid Occupancy map: 2D grid + resolution (m/cell).
Environment Area semantics Free text describing the area -+ e.g., “indoor", “outdoor", “kitchen", “a science museum".
TABLE I: Data stored per trajectory.
III-B Structure and storage
The raw dataset contains two main directories: one containing the trajectories and another one containing data about the raters and their ratings.
The trajectories directory contains JSON files of recorded trajectories in sub-directories named according to the source of the trajectory data. Each file contains one trajectory, and their relative file paths serve as trajectory identifiers. The data structure of these JSON files is presented in Table I.
The ratings directory contains a separate JSON file for each questionnaire completed, as described in III-A. In addition to the rater’s demographic information, each of these files store their ratings as a list of tuples: trajectory identifier (i.e., a relative path to the JSON file), context, score.
III-C Software layer
In the Git repository of the project [^2] https://github.com/SocNavData/SocNav3, we provide scripts to visualize and perform statistical analyses on the data, classes for data-loading, data augmentation and normalization, and code to deploy the surveying tool to acquire ratings.
Statistical analysis scripts: Used to compute the quadratic weighted Cohen’s kappa coefficient of the raters and to produce the consistency map in Fig. 4. More details are provided in Sec. IV-B.
Trajectory visualization tool: Unlike other types of datasets with simpler datatypes, such as plain text, images, or audio, trajectory data combines different variables, which makes it harder to work with and visualize. Visualization is crucial in this domain because the complexity of the data structures makes errors likely if the trajectories are not carefully compiled. Inconsistencies in coordinate or unit systems are common issues that can be picked up visually early on with the right tools. The visualization tool is shown in 2(a). A 3D visualization tool built on top of Webots [11] is also provided to generate video recordings.
(a) 2D custom representation.
(b) Webots-based 3D representation.
Figure 2: Trajectory visualization using a custom 2D representation (2(a)) and 3D simulator (Webots 2(b)).
Surveying tool: Running a ratings survey benefits from a more usable—and aesthetic—representation. The surveying tool shows video recordings along with a randomly generated context, asks raters to provide their scores, as shown in Fig. 3. The trajectories to use, the types of contexts, the maximum number of scores per rater, and the repeated control questions can all be configured.
Figure 3: Survey screenshot.
Data normalization, and augmentation: Normalization and augmentation help to include equivariances and invariances in machine learning models without embedding them into equations. Applying these transformations to robot trajectories is complex and error-prone, and so, as with visualization, tooling support is helpful. To avoid forcing users to duplicate data augmentation code, and to streamline and homogenize data augmentation and normalization, PyTorch data transforms implementing such features are provided.
Data normalization automatically reframes all poses into the goal frame of reference; this can steer machine learning models towards invariance w.r.t. the frame of reference.
The provided software implements three types of data augmentations. First, to enhance robustness to noise, poses can be augmented with additive Gaussian noise. Secondly, when goal orientations exceed a threshold larger or equal to $\pi$ rad, the orientation value is randomized. Finally, trajectory data can be randomly mirrored (left-right).
IV Description and Analysis of the Dataset Prototype
We provide a dataset to show how the proposed dataset structure can be exploited to generate a learning-based metric. We collected $4427$ trajectories, $182$ real and $4245$ simulated. We extended this initial set by varying the configuration of the walls in the scenarios of the compiled trajectories (i.e., removing some or all of the walls). This process resulted in a total of $7472$ trajectories. It is key to have both real and simulated data sources because they pose different challenges: real trajectories typically have lower variability than simulated ones because of the expense of recording trajectories in multiple scenarios, while simulated data have less realistic movements and noise. Additionally, due to safety and ethical reasons, dangerous robot trajectories can only be acquired in simulation. Furthermore, acquiring a large enough dataset to cover all possible situations with real data exclusively would be unrealistically expensive.
In this study, we kept cost bounded by limiting this proof-of-concept dataset in several ways, nevertheless ensuring good coverage within the selected domain. Firstly, the dataset is limited to indoor areas whose bounding box ranges between 6m6~m and 10m10~m wide. The navigation tasks covered in the dataset are all of the type “go-to”; other tasks such as following or guiding humans are left for future additions to the dataset repository. The objects considered in the dataset are limited to chairs, shelves, tables, lights, and plants. The maximum robot speed recorded is 2m/s2~m/s.
IV-A Trajectories recorded
Two sources were used to acquire real trajectories. A set of 109 trajectories was recorded at ARP laboratory (Aston University). A $8\times 7.4$ meters wide room was populated with tables, shelves, chairs, a TV, and from 1 to 6 humans. Another 73 trajectories were re-used from the SACSoN dataset [5]. The trajectories available in SACSoN were recorded in environments larger than the maximum area set for the proof-of-concept dataset, so they were cropped to comply with the range allowed within the restrictions for our dataset.
We included simulated trajectories from a range of simulators and algorithms, using both manual control and a learned RL agent on SocNavGym [6], HATEB [16], and HuNavSim [12], as well as manual control on Gazebo.
IV-B Data quality
Each survey included $15$ control questions to assess consistency between the participants. Additionally, five of these questions were presented twice to evaluate intra-rater consistency. In total, $49$ participants scored $6,481$ trajectories.
To ensure data quality, we first discarded ratings from participants who did not answer all control questions. For the $34$ remaining participants, we then applied a selection process that considers both intra-rater and inter-rater consistency using the quadratic weighted Cohen’s kappa coefficient. We define $\mu_{intraH}$ as a high-quality threshold and $\mu_{intraL}$ as a minimum-quality threshold for intra-rater consistency and $\mu_{inter}$ as a minimum threshold for inter-rater consistency. The selection consists of the following steps:
  1. 1.
    Compute the average score for each control question over the population of raters with an intra-rater consistency greater than $\mu_{intraH}$.
  2. 2.
    Discard raters with an intra-rater consistency lower than $\mu_{intraL}$ or an inter-rater consistency lower than $\mu_{inter}$ with respect to the average scores of the control questions.
Combining intra- and inter-rater agreements provides more solid results than a separate selection. Due to the limited number of repeated control questions, small variations in the answers of a rater could have a significant impact on their intra-rater agreement [19]. Thus, an initial selection based only on high intra-rater consistency could result in discarding too many raters. Using both coefficients, we can retain raters who align with the group opinion, even with minor inconsistencies in their responses. This approach improves the overall reliability and representativeness of the data by filtering out erratic and atypical ratings without being excessively restrictive.
Figure 4 shows the consistency map of the raters selected following the described procedure, considering values of $0.4$, $0.1$ and $0.2$ for $\mu_{intraH}$, $\mu_{intraL}$ and $\mu_{inter}$, respectively. The resulting set of 22 raters yields a final dataset of 4402 scored trajectories.
Figure 4: Consistency map of selected raters.
V Experimental Results
We propose a Gated Recurrent Unit-based neural network (GRU [3]) ALT metric model to demonstrate the feasibility of data-driven approaches and to serve as a baseline in future work. This section describes the data transformations applied, the architecture used, the training procedure, as well as quantitative and qualitative results.
V-A Data transformations
The raw trajectory data is augmented using mirroring and transformed into 1D vectors that serve as input to the GRU. The vectors ($x_{0},\dots,x_{t},\dots,x_{T}$) are the concatenation of trajectory features ($f_{t}$), metric-based features ($m_{t}$), and an ad-hoc context embedding ($c_{t}$):
[…]
The trajectory features, $f_{t}$, include pose, speed, and acceleration of the robot w.r.t. the goal frame of reference, goal thresholds (regarding position and orientation), general scenario information (flags indicating the presence of humans and the existence of walls), and data that allows contextualizing in time the current step in the whole trajectory (the current step ratio and a last step flag).
The metrics compiled in $m_{t}$ are the following:
  • A flag indicating whether the robot reaches the goal.
  • Distances to the nearest human, object, and wall.
  • Flags indicating whether the robot has collided with a human, an object, or a wall.
  • Number of humans within a radius of the robot ($0.4$, $0.6$, and 0.8m0.8~m).
  • Flags indicating social space intrusions, considering the same thresholds as in the previous point.
  • Features indicating potential danger to humans: minimum time to collision, maximum cost of fear, and maximum cost of panic as defined in [17].
  • Minimum distance from humans up to the current step.
  • Path efficiency ratio measured as the initial distance to the goal divided by the length of the trajectory up to the current step.
The contextual values in $c_{t}$ are computed from the free-form text describing the context of the task using queries to an LLM (Anthropic Claude). The prompt used was: “I will give a task description for a robot. I want you to reply with the percentile (a number from 0 to 100) that corresponds to <VARIABLE> in comparison with that of other tasks you could imagine. I don’t want an explanation, only the percentile. Take your time to think, but respond with a single integer from 0 to 100.”, where <VARIABLE> is substituted with each of the strings in Table II. The output of the LLM is later normalised to the range 0-1.
Contextual query
the urgency of the task
the importance of the task
the risk involved in the task
the distance the robot should keep to humans during the task
the distance the robot should keep to objects during the task
the speed with which the robot should move
the importance of comfort versus efficiency in the task
to what extent bumping into a human would be justified
to what extent bumping into an object would be justified
the importance of moving in a predictable way
TABLE II: Contextual queries used in the LLM prompts.
V-B Training procedure
After discarding the trajectories used as control questions, we create a 0.9/0.05/0.05 train/validation/test split and train a GRU-based neural network. The network consists of a GRU connected to an MLP. The GRU has 4 layers with 256-unit hidden layers, and the output of the last step of its sequence is fed into an MLP with a 256-unit hidden layer and a 1-unit output layer. The hidden layers use LeakyReLU activation with a $0.01$ negative slope, and the output layer uses sigmoid activation. The model is trained with a batch size of 32 and a learning rate of 1e-5, using MSE loss with the target of the raters’ scores. The training follows an early-stopping strategy with a patience of 20 epochs. Checkpoint saving is performed every time the loss of the validation set improves.
V-C Results
Training resulted in an MSE loss of $0.0469$ for the validation dataset, achieved in epoch $208$. The final test MSE loss is $\mathbf{0.0477}$, equivalent to a MAE of $\mathbf{0.162}$. The MSE and MAE of the model with respect to the mean of the control questions were $0.0062$ and $0.066$, respectively. To avoid over-representing these data points, control questions were not included in the dataset. Fig. 5 shows the mean scores for the control questions, along with their standard deviations, and the estimation the model made for each of them.
Figure 5: Plot for the control questions with their corresponding mean, standard deviation, and the estimations made by the model. The control questions are shown sorted according to their mean score.
To test the model from a qualitative perspective, we generated three simulated scenarios in which people are located frontally between the robot’s initial position and the goal position (Fig. 6(a)). The robot moves at different speeds following several trajectories (a total of $101$) that deviate a variable distance from the straight line to the goal. The top left image in Fig. 6(a) shows a subset of these trajectories as dotted lines. Additionally, four different contexts are considered for each scenario, each with distinct urgency, importance, and risk among other variables.
(a) Robot trajectories and scenarios of Figs. 6(b), 6(c), and 6(d).
(b) One static person.
(c) Three static persons.
(d) A person moving in the direction of the robot.
Figure 6: This figure shows the output of the learned metric for different trajectory variations, situations, contexts and speeds. In all scenarios, the robot and the goal are located in the same position, as indicated in Fig 6(a), in orange at the bottom and green at the top, respectively. In Figs. 6(b) and 6(d), there is a single human; in the first case static, in the second case approaching the location where the robot starts. In Fig.6(c), there are three humans, as depicted in Fig.6(a). To get to its goal the robot follows different trajectories, with varying degrees of divergence from the central line (top-left image of Fig. 6(a)). The contexts are as shown on the top-left of each figure: (lab) is for “A robot is working with lab samples. The samples contain a deadly virus”; (fire) is for “A restaurant robot is looking for a fire extinguisher, as it just detected a fire”; (office) is for “An office assistant robot keeps track of who is in the office today”, and (fragile) is for “A delivery robot is navigating in a hospital. It works with fragile objects”. Speed is color-coded:   0.20 m/s   0.40 m/s   0.80 m/s   1.60 m/s.
Figures 6(b), 6(c), and 6(d) show the metric model output for these three scenarios. The horizontal axis represents the maximum deviation of the trajectory from the straight line. The sign refers to the side, left or right, where the deviation occurs. The plots show the score assigned by the trained model to each trajectory.
As these figures show, the model learns to distinguish among contexts. The highest speed is penalized more in the contexts labeled as lab and fragile in Fig. 6, where the robot would generally be expected to move slowly due to the risk of the task. In these two contexts, the robot is also expected to maintain a wider distance from humans, which explains why trajectories closer to the human (i.e., with smaller deviations) receive lower scores than in the other contexts. In contrast, the fire context reflects a high-urgency situation: the lowest speed results in scores close to zero on all trajectories and scenarios, and closer proximity to humans is more acceptable.
In addition, in the office context, where the task is less critical, the lowest speed is less penalized and the scores decay more slowly than in the other contexts as the robot deviates from the straight line to the goal. In the scenario with three pedestrians (Fig 6(c)), this behavior is less evident. Nevertheless, it can be observed how, for the range of distances with the highest scores, the office context receives practically the same score, while the rest of the contexts are penalized as the deviation increases.
Finally, notable differences can be observed between the two scenarios with one person (Figs. 6(b) and 6(d)). When the person moves, the robot should move faster than when the person is static. This aligns with the expectation that the robot should actively maneuver in time to avoid the moving person safely. Additionally, deviating farther from the straight line is more penalized than in the static scenario, which can be explained by the fact that, once the robot has safely avoided the human, the distance from them is greater than in the stationary person scenario, as the person moves away from the robot.
VI Conclusions and Future Work
This paper described how All-encompassing Learned Task-wise metrics [4] can be learned following a data-driven approach in the context of social robot navigation. The experimental results support the idea that the approach is feasible; the quantitative results achieve MSE and MAE of $0.0477$ and $0.162$, respectively. The qualitative results show that the learned metric is sensitive to both the content of the scenarios and the context of the tasks.
The authors acknowledge these results could be enhanced with additional data and architectural improvements, which points to the need for further research in this area. More sophisticated architectures were not satisfactory in our experiments, arguably due to the size of the dataset.
Our envisioned future work spans two different areas: a) increasing the dataset size, and b) development of domain-specific deep learning architectures that can more efficiently leverage geometric relationships.